A company issues $15,000,000, 7.8%, 20-year bonds to yield 8% on January 1, 2020. Interest is paid on June 30 and December 31. The proceeds from the bonds are $14,703,108. What is interest expense for 2021, using straight-line amortization?
Answer #1
Discount on issue 296892 =15000000-14703108
Semi-annual discount amortization 7422.30 =296892/20*6/12
Semi-annual Interest payment 585000.00 =15000000*7.8%*6/12
Semi-annual Interest expense 592422.30
Interest expense for 2021 1184845 =592422.30*2